OpenStudy (anonymous):

can someone please check my work? Find a simple basic function as a right-end behavior model and a simple basic function as a left-end behavior model. y=3x^2 - 3sinx I got y = 3x^2; y=-3x^2

OpenStudy (anonymous):

\[F(x) = 0, x<0 \] \[F(x) = x^2 - 2x 0 - \le x \le2\] \[F(x) = 2 - 2>2 \]